MODERN ALGEBRA (MATH 210) PROBLEM SET 1

From the text: 2.1, 2.7, 2.15, 2.25, 2.27, 2.33, 2.43, 2.46, 2.52, 2.64 (iii), 2.65, 2.71. This set covers roughly up until quotient groups. If you get stuck, come see me, or talk with Pokman (the course assistant) during office

hours, or talk with other students.

The set is due Tuesday, October 15 at 3:30 pm in Pierre Albin’s mailbox (opposite the elevator on the first floor of Building 380). The score on your best ten problems will be counted.

Date: Tuesday, October 8, 2002.